OpenStudy (anonymous):

Use sum, differernce, product, or half-angle formulas to find the exact value of each expression tan 105 degrees I used the half angle formula, can someone check my answer and see if it's right Will FAN and MEDAL

OpenStudy (unknownunknown):

Go ahead and show your answer and working minisweet4.

OpenStudy (anonymous):

\[\sin \frac{ \theta }{ 2 }=+- \sqrt{1-5\pi/4}/4\] My final answer is sqrt 2+sqrt 2/2 @unknownunknown

OpenStudy (unknownunknown):

So is this to find the value of tan(105 degrees) ?

OpenStudy (anonymous):

You can try different formulas but this is the one I choose because my teacher said I did one thing wrong

OpenStudy (unknownunknown):

Alright, I still don't understand the question fully. In your original question you mention tan 105 degrees?

OpenStudy (anonymous):

Oh sorry I wrote out the wrong answer, hold on

OpenStudy (the_joker<3):

hmmm... what do you think?

OpenStudy (anonymous):

\[\tan \frac{ 210 }{ 2 }=\sqrt{1+\sqrt3/2} / \sqrt{1-\sqrt3/2}\] For the answer I got sqrt 5/1

OpenStudy (the_joker<3):